A random sample of 11 students provided the following age data: 25 19 20 21 19 23 22 19 33 21 24

Find the z-score for data value . (Round your answer to three decimal places.)
Answer #1

From the given data :

19,19,19,20,21,21,22,23,24,25,33

sample mean(xbar)=xi/n=246/11=22.3636

Sample standard deviation(S)=(xi-xbar)/(n-1)=[(19-22.3636)^2+...+(33-22.3636)^2]/(11-1))=4.081

Now Z score =(x,-xbar)/s